OpenStudy (anonymous):

Jamil always throws loose change into a pencil holder on his desk and takes it out every two weeks. This time it is all nickels and dimes. There are 7 times as many dimes as nickels, and the value of the dimes is $2.60 more than the value of the nickels. How many nickels and dimes does Jamil have? Choose one answer. a. 28 nickels and 4 dimes b. 3 nickels and 21 dimes c. 4 nickels and 28 dimes d. 5 nickels and 35 dimes Jamil always throws loose change into a pencil holder on his desk and takes it out every two weeks. This time it is all nickels and dimes. There are 7 times as many di

OpenStudy (anonymous):

35 dimes 5 nickels so answer D

OpenStudy (anonymous):

hey i solve this way round : I assume dime as d and nickel as n with their values as 0.1 and .05 of a dollar resp. then i formed 2 equations according to the question as d=7n and second one as \[0.1\times d =2.60 + 0.05 \times n\] then i plugged in the value of 1st equation in 2nd one as got n=4 and d=28 i.e. option C